# Streamlined Federated Unlearning: Unite as One to Be Highly Efficient

### v1

@@ Abstract
Recently, the enactment of “right to be forgotten” laws and regulations has imposed new privacy requirements on federated learning (FL). Researchers aim to remove the influence of certain data from the trained model without training from scratch through federated unlearning (FU). While current FU research has shown progress in enhancing unlearning efficiency, it often results in degraded model performance upon achieving the goal of data unlearning, necessitating additional steps to recover the performance of the unlearned model. Moreover, these approaches also suffer from many shortcomings such as high consumption of computational and storage resources. To this end, we propose a streamlined federated unlearning approach (SFU) aimed at effectively removing the influence of target data while preserving the model’s performance on the retained data without degradation. We design a practical multi-teacher system that achieves both target data influence removal and model performance preservation by guiding the unlearned model through several distinct teacher models. SFU is both computationally and storage-efficient, highly flexible, and generalizable. We conducted extensive experiments on both image and text benchmark datasets. The results demonstrate that SFU significantly improves time and communication efficiency compared to the benchmark retraining method and significantly outperforms existing state-of-the-art (SOTA) methods. Additionally, we verified the effectiveness of SFU using the backdoor attack.
@@ I Introduction
The open sharing of data is of significant importance in driving advancements across various fields, particularly in the rapidly evolving domain of artificial intelligence (AI). Building an effective machine learning model inevitably requires vast amounts of user data for training. However, during data circulation and usage, there exists a substantial risk of privacy leakage [1, 2]. Federated Learning (FL), as a distributed machine learning approach, was fundamentally designed to enable collaborative training while ensuring privacy protection [3]. Nonetheless, the recent enactment of regulations related to the ”right to be forgotten” has introduced new privacy protection requirements for FL [4, 5]. When users request the model to ”forget” a portion of data, simply deleting the targeted data from the database is insufficient, as the knowledge derived from this data is already embedded in the local models on each client through prior training.
@@
Evidently, the most straightforward approach to addressing this issue is to delete the target data from the database and retrain the model. However, this method incurs prohibitive computational costs. Moreover, in the context of FL, retraining may not even be feasible due to the privacy protection requirements of each participating entity. As a result, researchers aim to achieve data unlearning without retraining by utilizing a method known as Federated Unlearning (FU), which aims to accomplish the objective of forgetting targeted data and produce outcomes similar to those obtained through retraining. This approach seeks to substantially reduce the cost of unlearning while upholding data privacy and security.
@@
FU is designed to efficiently remove specific data and its influence from a trained model without retraining the entire model. However, FU faces unique challenges in the FL environment. Firstly, the presence of non-independently identically distributed (non-IID) data creates correlations and dependencies among the data, complicating the removal of specific data’s influence from the global model. Additionally, due to FL’s principle of data privacy protection, data cannot leave the local client, limiting the direct manipulation of raw data during FU and increasing the difficulty of designing effective unlearning mechanisms. Furthermore, compared to machine unlearning in centralized environments, FU must complete the unlearning task within a limited number of communication rounds, restricting the execution of complex operations.
@@
In recent years, FU research has made significant progress. As shown in Figure 1, the general workflow of current FU approaches typically involves two main steps: influence removal and performance recovery [6]. Influence removal entails applying unlearning techniques to eliminate the impact of specific target data from the model, aiming to make the behavior of the unlearned model as though it had never encountered the target data. This step has seen a variety of approaches. For instance, gradient manipulation techniques reduce the influence of the target data by adding noise or adjusting gradient directions[6, 34, 70, 80]. However, these approaches require substantial computational resources to determine appropriate noise levels and gradient directions, and the added noise can degrade the model’s performance on retained data. Other influence removal approaches include leveraging historical information [7, 8, 9, 10, 11], loss function approximations[12, 13], knowledge distillation[14, 10, 15] and so on. However, these techniques often incur additional storage costs, inaccuracies in estimating data influence, and other limitations. Most critically, the influence removal step in the majority of FU approaches tends to degrade the model’s performance on non-target data. This necessitates an additional performance recovery step, typically carried out through post-training techniques. Post-training [7, 16, 17] enables the model to relearn essential patterns after influence removal but adds both computation time and resource costs. Performance recovery can also be achieved through fine-tuning [18, 13, 19, 20], regularization [15, 14, 21], and similar techniques, which unavoidably increase the computational burden.
@@
To further improve unlearning efficiency, we considered can we streamline the steps of FU by combining influence removal and performance recovery into a single step? To this end, we propose a streamlined federated unlearning method called SFU. Inspired by knowledge distillation, SFU introduces a multi-teacher system, where multiple distinct teacher models guide the unlearning model (student model) to remove the influence of target data while minimizing performance degradation on the retained data. These teacher models require no additional training, thus avoiding extra computational costs. Furthermore, SFU does not rely on storing historical information or require additional storage resources. SFU demonstrates excellent versatility, as it can be applied to both traditional simple machine learning models and complex deep learning models. Given the constraints of FL environments, SFU operates without the need for global data access. Instead, each client performs unlearning on its own data, with no restrictions on data distribution. More importantly, SFU offers high flexibility, as it does not interrupt the existing FL process. It can seamlessly integrate into any FL training phase when unlearning is required, and after unlearning, the original FL training can resume. For categorical data unlearning, we conducted experiments on image and text benchmark datasets. Results show that SFU significantly improves both time and communication efficiency compared to baseline retraining methods, and far surpasses existing state-of-the-art (SOTA) approaches. Additionally, we validated its effectiveness using backdoor attacks.
@@
  1. 1)
    We propose a streamlined federated unlearning approach, SFU, which combines the two steps of unlearning’s influence removal and performance recovery into one, dramatically improving unlearning efficiency.
  2. 2)
@@
  3. 3)
    SFU is perfectly adapted to FL environments and does not require global data access. Instead, individual clients train on their own data without restrictions on data distribution. Additionally, it offers great flexibility, seamlessly transitioning to any training phase in FL.
  4. 4)
    Extensive experimental results on image and text datasets show that SFU significantly improves time efficiency and communication efficiency compared to benchmark retraining methods and significantly outperforms existing state-of-the-art (SOTA) methods. Furthermore, we utilize a backdoor attack to verify its effectiveness.
The rest of this paper is organized as follows. In Section II, we reviewed the relevant literature on federated unlearning. In Section III, we introduced the necessary background knowledge for SFU. Section IV provides a detailed description of the proposed SFU method and its specific implementation. Section V presents the experimental results, and Section VI concludes the paper.
@@ II-A Federated unlearning
Due to the infeasibility of retraining in Federated Learning (FL), current research on Federated Unlearning (FU) has explored various directions. FedRecover[7] tracks the updates from each client during every training round and removes the updates of the target clients during unlearning, using only the retained clients’ updates to estimate the unlearned model. Similarly, MetaFul[9] removes the influence of target data by subtracting the historical updates of the target data from the global model. FedEraser[22] builds on this by employing direction calibration, storing only the most recent updates instead of all historical updates. Other similar methods, such as[23, 24, 25], rely on historical updates to improve unlearning efficiency to a small extent, but these approaches require large storage resources to maintain historical data. FFMU[26] improves robustness to unlearning by adding Gaussian noise to smooth the local model gradients and performs unlearning of the global model during aggregation using methods like the Nemytskii operator. URKL[27] trains noise input to perturb the target model’s gradient updates, facilitating unlearning. In FedFilter[28], the server generates random reverse gradients and performs training using stochastic gradient descent (SGD) to remove the influence of target data. Verifi[29] achieves unlearning of target clients by shrinking the gradients of the target client and amplifying those of the remaining clients. FUMD[30] uses a momentum decay strategy to gradually reduce the model’s ability to discriminate the target data, effectively removing its influence. FUCDP[31] uses TF-IDF to calculate the correlation scores between channels and categories and prunes the channels most similar to the target category to complete unlearning. These methods can generally be classified as gradient-based unlearning techniques, which tend to be computationally efficient but are challenging to apply without negatively affecting the overall model performance. Liu et al.[13] and Jin et al.[12] adjust model updates by approximating higher-order derivatives of the loss function, which theoretically offers more refined control but faces significant computational complexity challenges and may lead to performance degradation. Reverse training methods, which maximize the loss using gradient ascent instead of descent, are intuitive and effective but significantly reduce model performance and thus require additional performance recovery methods to mitigate the negative impact[32, 33, 34].
@@ II-B Knowledge distillation in FU
While some existing FU studies have employed knowledge distillation to develop unlearning methodologies, these approaches have been found to have inherent limitations. FUAF[14] utilizes a teacher model to generate pseudo-labels with original features for influence removal. However, it still requires elastic weight consolidation for performance recovery. Wu et al.[10] achieve influence removal of backdoor information by using historical client updates, but this leads to performance degradation. To address this, they employ the original model as a teacher model to guide the unlearned model through the performance recovery step. QuickDrop[15] generates a representative small synthetic dataset through dataset distillation and then uses this distilled dataset to perform Stochastic Gradient Ascent (SGA) for effective influence removal of target data. However, it still requires the use of regular Stochastic Gradient Descent (SGD) on the distilled data for performance recovery. These methods address either the influence removal or performance recovery step through the teacher-student model, but they still face limitations in terms of unlearning efficiency. The necessity of separate steps for influence removal and performance recovery, along with additional computational overhead, results in suboptimal unlearning efficiency.
@@ II-C Discussion
Overall, the vast majority of existing FU methods lead to a decline in model performance when removing the influence of target data, requiring an additional performance recovery step. Although FU methods based on historical information may avoid performance degradation, their efficiency improvements are minimal, and they require substantial storage resources. FedRecovery[11] introduces noise and utilizes differential privacy to make the unlearned model indistinguishable from a retrained model. While they claim to eliminate the performance recovery step, the performance of the unlearned model still lags behind that of the retrained model, a gap that cannot be ignored. To better address these issues, we propose a streamlined FU method called SFU. SFU is based on a specific multi-teacher system, which not only achieves influence removal but also preserves model performance simultaneously. This effectively combines the influence removal and performance recovery steps into one, significantly enhancing unlearning efficiency.

### v2

@@ Abstract
Recently, the enactment of “right to be forgotten” laws and regulations has imposed new privacy requirements on federated learning (FL). Researchers aim to remove the influence of certain data from the trained model without training from scratch through federated unlearning (FU). While current FU research has shown progress in enhancing unlearning efficiency, it often results in degraded model performance upon achieving the goal of data unlearning, necessitating additional steps to recover the performance of the unlearned model. Moreover, these approaches also suffer from many shortcomings such as high consumption of computational and storage resources. To this end, we propose a streamlined federated unlearning approach (SFU) aimed at effectively removing the influence of the target data while preserving the model performance on the retained data without degradation. We design a practical multi-teacher system that achieves both target data influence removal and model performance preservation by guiding the unlearned model through several distinct teacher models. SFU is both computationally and storage-efficient, highly flexible, and generalizable. We conduct extensive experiments on both image and text benchmark datasets. The results demonstrate that SFU significantly improves time and communication efficiency compared to the benchmark retraining method and significantly outperforms existing SOTA methods. Additionally, we verify the effectiveness of SFU using the backdoor attack.
@@ I Introduction
The open sharing of data is of significant importance in driving advancements across various fields, particularly in the rapidly evolving domain of artificial intelligence. Building an effective machine learning model inevitably requires vast amounts of user data for training. However, during data circulation and usage, there exists a substantial risk of privacy leakage [1, 2]. Federated learning (FL), as a distributed machine learning approach, was fundamentally designed to enable collaborative training while ensuring privacy protection [3]. Nonetheless, the recent enactment of regulations related to the “right to be forgotten” has introduced new privacy protection requirements for FL [4, 5]. When users request the model to “forget” a portion of data, simply deleting the targeted data from the dataset is insufficient, as the knowledge derived from this data is already embedded in the local models on each client through prior training.
@@
Evidently, the most straightforward approach to addressing this issue is to delete the target data from the dataset and retrain the model. However, this method incurs prohibitive computational costs. Moreover, in the context of FL, retraining may not even be feasible due to the privacy protection requirements of each participating entity. As a result, researchers aim to achieve data unlearning without retraining by utilizing a method known as federated unlearning (FU) [6], which aims to accomplish the objective of forgetting targeted data and producing outcomes similar to those obtained through retraining. This approach seeks to reduce the cost of unlearning substantially while upholding data privacy and security.
@@
FU is designed to efficiently remove specific data and its influence from a trained model without retraining the entire model. In contrast to machine unlearning (MU) in centralized environments [7, 8, 9, 10], FU faces unique challenges in FL environments. This makes many MU methods not directly translatable to the ideas of FU methods, thereby exacerbating the design complexity. Firstly, the presence of non-independently identically distributed (non-iid) data creates correlations and dependencies among the data, complicating the influence removal of specific data from the global model. Additionally, due to the principle of data privacy protection in FL, data cannot leave the local client, limiting the direct manipulation of raw data during FU and increasing the difficulty of designing effective unlearning mechanisms. Furthermore, compared to MU, FU must complete the unlearning task within a limited number of communication rounds, restricting the execution of complex operations.
@@
In recent years, FU research has made significant progress. As shown in Figure 1, the general workflow of current FU approaches typically involves two main steps: influence removal and performance recovery [11]. Influence removal entails applying unlearning techniques to eliminate the impact of specific target data from the model, aiming to make the behavior of the unlearned model as though it had never encountered the target data. This step has seen a variety of approaches. For instance, gradient manipulation techniques reduce the influence of the target data by adding noise or adjusting gradient directions[6, 34, 70, 80]. However, these approaches require substantial computational resources to determine appropriate noise levels and gradient directions, and the added noise can degrade the model performance on retained data. Other influence removal approaches include leveraging historical information [12, 13, 14, 15, 16], loss function approximations [17, 18], knowledge distillation [19, 15, 20] and so on. However, these techniques often incur additional storage costs, inaccuracies in estimating data influence, and other limitations. Most critically, the influence removal step in the majority of FU approaches tends to degrade the model performance on non-target data. This necessitates an additional performance recovery step, typically carried out through post-training techniques. Post-training [12, 21, 22] enables the model to relearn essential patterns after influence removal but adds both computation time and resource costs. Performance recovery can also be achieved through fine-tuning [23, 18, 24, 25], regularization [20, 19, 26], and similar techniques, which unavoidably increase the computational burden.
@@
To further improve unlearning efficiency, we considered can we streamline the steps of FU by combining influence removal and performance recovery into a single step? To this end, we propose a streamlined federated unlearning method called SFU. Inspired by knowledge distillation, SFU introduces a multi-teacher system, where multiple distinct teacher models guide the unlearning model (student model) to remove the influence of target data while minimizing performance degradation on the retained data. These teacher models require no additional training, thus avoiding extra computational costs. Furthermore, SFU does not rely on storing historical information or require additional storage resources. SFU demonstrates excellent versatility, as it can be applied to both traditional simple machine learning models and complex deep learning models. Given the constraints of FL environments, SFU operates without the need for global data access. Instead, each client performs unlearning on its own data, with no restrictions on data distribution. More importantly, SFU offers high flexibility as it does not interrupt the existing FL process. It can seamlessly integrate into any FL training phase when unlearning is required, and after unlearning, the original FL training can resume. For categorical data unlearning, we conducted experiments on image and text benchmark datasets. Results show that SFU significantly improves both time and communication efficiency compared to baseline retraining methods, and far surpasses existing state-of-the-art (SOTA) approaches. Additionally, we implant backdoors on the forgotten data and verify the effectiveness of SFU by comparing the success rate of backdoor attacks before and after unlearning.
@@
  1. 1.
    We propose a Streamlined Federated Unlearning approach called SFU, which combines the two steps of influence removal and performance recovery into one, dramatically improving unlearning efficiency.
  2. 2.
@@
  3. 3.
    SFU can be well adapted to FL environments and does not require global data access. Instead, individual clients train on their own data without restrictions on data distribution. Additionally, it offers great flexibility, seamlessly transitioning to any training phase in FL.
  4. 4.
    Extensive experimental results on image and text datasets show that SFU significantly improves time efficiency and communication efficiency compared to benchmark retraining methods and significantly outperforms existing SOTA methods. Furthermore, we verify the effectiveness of SFU by comparing the success rate of backdoor attacks on embedded backdoors in forgotten data before and after unlearning.
The rest of this paper is organized as follows. In Section II, we review the relevant literature on federated unlearning. In Section III, we introduce the necessary background knowledge for SFU. Section IV provides a detailed description of the proposed SFU method and its specific implementation. Section V demonstrates the theoretical analysis of SFU. Section VI presents the experimental results, and Section VII concludes the paper.
@@ II-A Federated Unlearning
Due to the infeasibility of retraining in FL, current research on FU has explored various directions. FedRecover [12] tracks the updates from each client during every training round and removes the updates of the target clients during unlearning, using only the retained clients’ updates to estimate the unlearned model. Similarly, MetaFul [14] removes the influence of target data by subtracting the historical updates of the target data from the global model. FedEraser [27] builds on this by employing direction calibration, storing only the most recent updates instead of all historical updates. Other similar methods, such as [28, 29, 30], rely on historical updates to improve unlearning efficiency to a small extent, but these approaches require large storage resources to maintain historical data. FFMU [31] improves robustness to unlearning by adding Gaussian noise to smooth the local model gradients and performs unlearning of the global model during aggregation using methods like the Nemytskii operator. URKL [32] trains noise input to perturb the gradient updates of the target model, facilitating unlearning. In FedFilter [33], the server generates random reverse gradients and performs training using stochastic gradient descent (SGD) to remove the influence of target data. Verifi [34] achieves unlearning of target clients by shrinking the gradients of the target client and amplifying those of the remaining clients. FUMD [35] uses a momentum decay strategy to gradually reduce the model ability to discriminate the target data, effectively removing its influence. FUCDP [36] uses TF-IDF to calculate the correlation scores between channels and categories and prunes the channels most similar to the target category to complete unlearning. These methods can generally be classified as gradient-based unlearning techniques, which tend to be computationally efficient but are challenging to apply without negatively affecting the overall model performance. Liu et al. [18] and Jin et al. [17] adjust model updates by approximating higher-order derivatives of the loss function, which theoretically offers more refined control but faces significant computational complexity challenges and may lead to performance degradation. Reverse training methods, which maximize the loss using gradient ascent instead of descent, are intuitive and effective but significantly reduce model performance and thus require additional performance recovery methods to mitigate the negative impact [37, 38, 39].
@@ II-B Knowledge Distillation in FU
While some existing FU studies have employed knowledge distillation to develop unlearning methodologies, these approaches have been found to have inherent limitations. FUAF [19] utilizes a teacher model to generate pseudo-labels with original features for influence removal. However, it still requires elastic weight consolidation for performance recovery. Wu et al. [15] achieve influence removal of backdoor information by using historical client updates, but this leads to performance degradation. To address this, they employ the original model as a teacher model to guide the unlearned model through the performance recovery step. QuickDrop [20] generates a representative small synthetic dataset through dataset distillation and then uses this distilled dataset to perform stochastic gradient ascent for effective influence removal of target data. However, it still requires the use of regular stochastic gradient descent (SGD) on the distilled data for performance recovery. These methods address either the influence removal or performance recovery step through the teacher-student model, but they still face limitations in terms of unlearning efficiency. The necessity of separate steps for influence removal and performance recovery, along with additional computational overhead, results in suboptimal unlearning efficiency.
@@ II-C Discussion
As shown in table I, the vast majority of existing FU methods lead to a decline in model performance when removing the influence of target data, requiring an additional performance recovery step. Although FU methods based on historical information may avoid performance degradation, their efficiency improvements are minimal, and they require substantial storage resources. FUR [16] introduces noise and utilizes differential privacy to make the unlearned model indistinguishable from a retrained model. While they claim to eliminate the performance recovery step, the performance of the unlearned model still lags behind that of the retrained model, a gap that cannot be ignored. FedU [40] leverages influence functions to estimate and remove the influence of deleted samples from model parameters for unlearning. To counteract performance degradation, they retrain the model on retained data and introduce an adaptive optimization method to balance forgetting and performance retention. However, this balance limits efficiency improvements to a certain extent. Moreover, FedU is specifically designed for unlearning a small number of samples and cannot be easily extended to other unlearning objectives, such as class-level unlearning. To better address these issues, we propose a streamlined FU method called SFU. SFU is based on a specific multi-teacher system, which not only achieves influence removal but also preserves model performance simultaneously. This effectively combines the influence removal and performance recovery steps into one, significantly enhancing unlearning efficiency.
TABLE I: Summary and comparison
Approaches
Influence
removal
High
efficiency
Performance
maintenance
No fine-tuning
required
Storing historical
information
[12, 14, 27, 28, 29, 30]
$\checkmark$ $\times$ $\checkmark$ $\checkmark$
Gradient-based
[31, 32, 33, 34, 35, 36]
$\checkmark$ $\checkmark$ $\times$($\downarrow$) $\times$
Loss function
approximation
[18, 17]
$\checkmark$ $\checkmark$ $\times$($\downarrow$) $\times$
Reverse training
[37, 38, 39]
$\checkmark$ $\times$ $\times$($\downarrow$) $\times$
Knowledge distillation
[19, 15, 20]
$\checkmark$ $\checkmark$ $\times$($\downarrow$) $\times$
Differential privacy
[16]
$\checkmark$ $\checkmark$ $\times$($\downarrow$) $\times$
SFU $\checkmark$ $\checkmark$(highest) $\checkmark$ $\checkmark$
@@ III Preliminary
In this section, we introduce the background knowledge and fundamental concepts related to federated learning involved in SFU. Basic notations and their meanings used in this paper can be found in Table II.
TABLE II: Basic Notations
Notations Descriptions
$KL$ Kullback-Leibler divergence
$M^{P}$ The performance preservation teacher model
$M^{F}$ The forgetting teacher model
$M_{i}^{L}$ The label-based preservation teacher model
$M_{i}^{U}$ Unlearned model on client $i$
$M^{U}$ Global unlearned Model
$x_{i}$ Total data on client $i$
$x_{i}^{f}$, $x_{i}^{r}$ Data that needs to be forgotten/retained on client $i$
$\alpha$ Hyperparameters
$num(x)$ Number of samples $x$
$one\_hot$ One-hot encoding
$y(x)$ Labels of samples $x$
FL is a decentralized machine learning approach where multiple clients collaboratively train a global model without sharing their private data. Federated averaging (FedAvg) as a common aggregation method is used in various scenarios [3]. Each client holds its own local dataset $x_{i}$, and the goal is to learn a global model parameter $M^{*}$ that minimizes the overall loss across all clients. In each iteration, the server sends the current global model $M_{t}$ to all clients. Each client then performs local training on its dataset using SGD for several epochs, producing an updated model $M_{i,t+1}$. The server aggregates the model updates from all clients using a weighted average:
$M_{t+1}=\sum_{i=1}^{I}w_{i}M_{i,t+1},w_{i}=\frac{num(x_{i})}{\sum_{i=1}^{I}num(x_{i})},\vspace{5pt}$ (1)
where $w_{i}$ is the weight assigned to each client based on the size of its dataset, with $num(x_{i})$ being the number of samples $x_{i}$ in the dataset of the client $i$. The objective is to minimize the global loss function:
$M^{*}=\arg\min_{M}\frac{1}{I}\sum_{i=1}^{I}w_{i}L_{i}(M).\vspace{5pt}$ (2)
IV Our Approach
In this section, we first introduce the rationale for the teacher-student model and then introduce the design philosophy of each teacher model in the multi-teacher system used in SFU. Then, we provide a detailed description of the overall structure of the system, and finally, we present the specific implementation of SFU.
IV-A Teacher-student Model
The concept of the teacher-student model originated from knowledge distillation and is widely used in FL [41, 42]. The principle of a general teacher-student model is illustrated below:
IV-A1 Step 1: Knowledge Extraction
In the initial phase of the training process, the knowledge base is extracted from the teacher model. Consider a scenario where the teacher model has an output $q$, and the student model has an output $p$. These outputs are typically represented as probability distributions. The raw output of each model is converted into probabilities using the softmax function:
$q=softmax(Z_{t}/T),$ (3)
$p=softmax(Z_{s}/T).\vspace{5pt}$ (4)
IV-A2 Step 2: Knowledge Transfer
The student model then needs to receive knowledge from the teacher model, thereby reducing the difference between them. In knowledge distillation, the difference between the teacher model and the student model is usually described by KL divergence (Kullback-Leibler divergence). $KL(q||p)$ represents the loss of information when the true distribution $q$ of the teacher model is approximated by the distribution $p$ of the student model. In the teacher-student model, the goal is to minimize this difference so that the student model increasingly approximates the teacher model. Thus, the training objective is to minimize:
$KL(q\|p)=\sum_{k}q_{k}log\frac{q_{k}}{p_{k}},$ (5)
where $k$ represents the class index, and $q_{k}$ and $p_{k}$ are the predicted probabilities of the teacher model and the student model for the $k$-th class, respectively.
When multiple teachers are involved, it can be understood that the student model learns different knowledge from multiple different teacher models simultaneously. The training objective in this scenario is to minimize:
$KL(q_{1}\|p)+\dots+KL(q_{j}\|p)$ (6)
$=\sum_{k}q_{1,k}log\frac{q_{1,k}}{p_{k}}+\dots+\sum_{k}q_{j,k}log\frac{q_{j,k}}{p_{k}},$
where $q_{1,k},\dots,q_{j,k}$ and $p_{k}$ are the predicted probabilities of the $j$ teacher models and the student model for the $k$-th class, respectively. By minimizing the above loss function, the student model can be made to approach the outputs of multiple teacher models, thus integrating the knowledge from multiple teacher models.